# Cosmological simulation of radio synchrotron bridge between pre-merging galaxy clusters

### 10 Mar 2026 v1

@@ Abstract
Methods. We perform a cosmological magneto-hydrodynamical (MHD) simulation using the Enzo code. We improved a run-time Lagrangian tracer method implemented in Enzo, and follow the trajectories of baryonic matter using $N=\mathcal{O}(10^{7})$ tracer particles. In post-processing, we conduct a parallel computation of the Fokker–Planck (FP) equation for all tracers, with cooling and reacceleration efficiencies evaluated from the local quantities recorded along each tracer trajectory.
Results. Our simulation generate a Mpc-sized radio bridge in the early stage of the cluster merger. Within a reasonable parameter range, the reacceleration model produces a broad variety of spectra. In our fiducial model, the simulated bridge matches several properties of the one found between Abell 399 and Abell 401, such as its spectral shape, intensity profile, and pixel-by-pixel correlation between radio and X-ray intensities.
@@ 1 Introduction
In the past decades, radio observations have revealed a variety of diffuse synchrotron emission in GCs, such as radio halos and cluster radio shocks (van Weeren et al. 2019). The unprecedented sensitivity of recent radio facilities, such as Low Frequency Array (LOFAR) and MeerKAT, has enabled the detection of diffuse emission on larger scales (e.g., Rajpurohit et al. 2021; Botteon et al. 2022; Cuciti et al. 2022; van Weeren et al. 2024; Rajpurohit et al. 2025). Govoni et al. (2019) reported the first discovery of a “radio bridge” at 144 MHz in A399–A401, which is a diffuse emission covering the region between a pair of merging GCs.
@@
Another radio bridge has been found in Abell 1758 system (Botteon et al. 2020), which consists of two massive galaxy clusters. However, non-detections in other similar systems indicate that such bridge-like emission is not ubiquitous in the filamentary regions between clusters (e.g., Botteon et al. 2019; Brüggen et al. 2021; Pignataro et al. 2024a). The detection of radio bridges suggests that magnetic fields and cosmic-ray electrons (CREs) can permeate into the cosmic web beyond cluster virial radii. In contrast, the non-detections indicate that specific physical conditions are required to generate detectable radio bridge, or that the phenomenon is transient on cosmological timescales.
@@
The mechanisms responsible for magnetic field amplification and particle acceleration in inter-cluster bridge remain uncertain. The CREs in the bridge region may be injected through the feedback processes associated with active galactic nuclei (AGNs) or star-forming galaxies (e.g., Völk et al. 1996; Berezinsky et al. 1997). However, once injected, the CREs are subjected to severe energy losses owing to synchrotron and inverse-Compton radiation, as well as Coulomb collisions. As a result, the lifetime of radio-emitting CREs ($\sim 100$ Myr) is too short for them to propagate over megaparsec scales and to produce smooth, extended bridge emission. The same problem holds for giant radio halos in GCs and is commonly referred to as the slow-diffusion problem Brunetti and Jones (2014). Reacceleration of CREs is therefore often invoked as a solution to this problem in the cluster environments (e.g., Brunetti et al. 2001; Peterson et al. 2001; Fujita et al. 2003).
@@
Numerical simulations have shown that the intra-cluster medium (ICM) and the inter-cluster bridges are permeated by shocks and turbulence driven by the continuous matter accretion and episodic merger events of dark matter halos (e.g., Ryu et al. 2003; Pfrommer et al. 2006; Nelson et al. 2014; Steinwandel et al. 2024). Stochastic Fermi acceleration at these shocks or within turbulent flows may not be efficient enough to accelerate CREs directly from the thermal pool, but it can counteract radiative energy loss of the pre-existing CREs (e.g., Brunetti and Jones 2014). Secondary electrons generated through the hadronic interaction of CR protons can also contribute to diffuse synchrotron emission (e.g., Dennison 1980; Blasi and Colafrancesco 1999). However, in low-density gas regions such as cluster outskirts, the contribution from secondary electrons is expected to be smaller than in the central regions of clusters.
@@
Several previous works attempted to model energy-resolved CR distributions in GCs and LSS by combining cosmological MHD simulations with FP solvers. While earlier studies relied on MHD snapshots in post-processing (e.g., Donnert and Brunetti 2014; Pinzke et al. 2017; Winner et al. 2019; Smolinski et al. 2023; Beduzzi et al. 2024), more recent frameworks implement on-the-fly FP solvers under approximations that reduce the number of energy or momentum bins (e.g., Hopkins et al. 2022; Böss et al. 2023). We extend these approaches by incorporating the tracer-particle code Crater (Wittor et al. 2017) into the ENZO framework, and performing FP simulations in post-processing. The tracer particles generated and propagated during the runtime of Enzo better represent the baryonic flow compared to post-processing approaches such as Crater. We apply the FP solver developed in Nishiwaki et al. (2024); Nishiwaki and Asano (2025). Tests of the tracer and FP methods are presented in Appendix A.
@@
Figure 1: Evolution of the projected gas density map along Z axis in our Enzo simulation. The maps of zoomed-in region centered on the inter-cluster filament are shown. The white square shows the (1.2 Mpc)3 box region where a radio bridge forms (Sect. 3.1).
@@ 2 Method
In Sect. 2.1, we present the set-up of the Enzo simulation of the merging clusters. During the run of the simulation, $N\approx 1.5\times 10^{7}$ tracer particles are generated to sample the baryonic fluid elements. In Sect. 2.2, we explain the methods for tracer particle generation and propagation in the Enzo simulation. The strength of solenoidal turbulence around each tracer particle is quantified using the local vorticity (Sect. 2.3). The magnetic field in our Enzo simulation may lead to an underestimation of magnetic-field amplification by the turbulent dynamo due to the limited spatial resolution as well as partially due to the used Dedner hyperbolic MHD method. For this reason, we estimate a dynamo-amplified field in post-processing (Sect. 2.4) and adopt the larger of the simulated and modeled values. In Sect. 2.5, we introduce the FP equation adopted in this work and summarize the model parameters.
@@ 2.1 Enzo simulation
We simulate a pair of merging galaxy clusters by running the cosmological MHD code Enzo (Bryan et al. 2014). We perform a new re-simulation of the same system studied in Govoni et al. (2019) to make use of our developed tracer particle method on board Enzo. The simulation box is centered on the inter-cluster filament between merging clusters, resembling the A399–A401 system at low redshift. The root grid covers a comoving volume of $(260\penalty 10000\ \mathrm{Mpc})^{3}$, resolved with $220^{3}$ cells. Following Govoni et al. (2019), we employ the MHD solver with Dedner divergence cleaning (Dedner et al. 2002). We set an uniform comoving magnetic field with components $(B_{x},B_{y},B_{z})=(10^{-10},\,10^{-10},\,10^{-10})\penalty 10000\ \mathrm{G}$ at the initial redshift $z=30$ of the simulation. The radiative cooling of the baryonic gas is not included in this simulation.
We employ the adaptive mesh refinement (AMR) in the innermost (30 Mpc)3 region centered on the inter-cluster filament. Refined grids are introduced in the regions where the mass density of the baryonic gas or dark matter particles exceeds $0.1\rho_{\rm m}$, where $\rho_{\rm m}$ denotes the mean mass density of the Universe. The cell width is reduced by a factor of two at each refinement level. The maximum AMR level is eight, which corresponds to 4.2 comoving kpc per cell. From $z=2$, we decrease the refinement threshold to $0.01\rho_{\rm m}$ and introduce a super-Lagrangian level dependence of the refinement threshold with the exponent of 0.8-0.8. With this prescription, the refinement threshold decreases by a factor of $2^{-0.8}\simeq 0.57$ at each successive AMR level. This aggressive criteria ensures that most of the bridge volume with $\rho_{\rm gas}\gtrsim 10^{-28}$ g cm-3 is resolved with the resolution of AMR level seven (8.4 comoving kpc) in $z\lesssim 0.15$.
@@ 2.2 Lagrangian tracers
In many previous studies, the generation and propagation of Lagrangian tracer particles have been done in post-processing (e.g., Vazza et al. 2010; Wittor et al. 2017; Beduzzi et al. 2023; Smolinski et al. 2023). In this approach, particle positions are updated according to the local fluid velocity obtained by interpolating grid data from two consecutive simulation snapshots. However, because time interval between snapshots is generally much larger than the hydrodynamical or MHD time step at high AMR levels, this method can lead to systematic over- or undershooting of tracer particle positions.
@@
We iterate over all grid cells within the $(20\penalty 10000\ \mathrm{Mpc})^{3}$ region and evaluate the baryonic mass of each cell, $M_{\rm cell}$. If $M_{\rm cell}$ lies within the range $M_{\rm trc}^{\rm min},\,M_{\rm trc}^{\rm max}]$, the cell mass is represented by a single tracer particle with $M_{\rm trc}=M_{\rm cell}$. When $M_{\rm cell}>M_{\rm trc}^{\rm max}$, which typically occur at coarse resolution, the mass is represented by $\left\lfloor M_{\rm cell}/M_{\rm trc}^{\rm max}\right\rfloor$ tracer particles with $M_{\rm trc}=M_{\rm trc}^{\rm max}$, with one additional particle accounting for the residual mass. When $M_{\rm cell}<M_{\rm trc}^{\rm min}$, the cell mass is sampled probabilistically by generating a tracer particle with mass $M_{\rm trc}^{\rm min}$ with a probability of $M_{\rm cell}/M_{\rm trc}^{\rm min}$.
@@ 2.3 Measuring the local turbulent field
Solenoidal turbulence may play an important role in both magnetic-field amplification by the turbulent dynamo and CR reacceleration in the ICM (e.g., Cho and Vishniac 2000; Brunetti and Lazarian 2016; Vazza et al. 2018; Nishiwaki et al. 2024). Although the compressive component of turbulence may also contribute to CR acceleration (e.g., Brunetti and Lazarian 2007), we focus exclusively on the solenoidal mode in this work and defer a detailed comparison of between the two modes to future studies. The local solenoidal turbulent velocity is estimated from the vorticity using the relation $\delta v(l)\approx l|\nabla\times\bm{v}|$, where $l$ denotes the spatial scale at which the vorticity is evaluated. We calculate $\nabla\times\bm{v}$ using the central-difference scheme applied to the velocity field on a stencil of $3^{3}$ cells surrounding the nearest grid cell of each tracer particle. With this approach, tracer particles probe turbulent motions on different scales, depending on the resolution of their host AMR grid. For the calculation of the dynamo-amplified magnetic field and the CR acceleration efficiency in the following sections, we normalize the turbulent velocity $\delta v$ to a reference scale of $l=L=150$ comoving kpc, assuming a Kolmogorov scaling relation $\delta v(l)\propto l^{1/3}$.
@@ 2.4 Magnetic-field amplification through turbulent dynamo
The turbulent dynamo is one of the most plausible mechanisms for the amplification of magnetic fields in galaxy clusters. Gas in the inter-cluster bridge has densities and temperatures similar to those of the cluster outskirts ($n_{e}\sim 10^{-4}$ cm-3 and $T\sim 3$ keV), so constitutes a weakly-collisional high-beta plasma. Under these conditions, microphysical instabilities can suppress the effective viscosity and allow super-Alfvénic turbulence to cascade down to small scales (e.g., Schekochihin and Cowley 2006; Brunetti and Lazarian 2007; Beresnyak and Miniati 2016; Brunetti and Jones 2014; Kunz et al. 2022). However, resolving the turbulent dynamo directly in numerical simulations is computationally very demanding. The characteristic scale of this is the so-called the Alfvén scale,
@@ 2.5.3 Initial cosmic-ray distribution
Figure 5: From top to bottom, we show time evolution of the X-ray peak distance between the clusters, baryon mass, solenoidal turbulent energy flux, and the 140 MHz radio power, and the typical equilibrium energy of CREs. The latter four quantities are calculated for the (1.2 Mpc)3 box region extracted from the inter-cluster filament (Fig. 1).
@@ 3.1 Formation of a radio bridge
To study the properties of the radio bridge, we extract a box region in the inter-cluster filament with a volume of $(1.2\penalty 10000\ {\rm Mpc})^{3}$. The region is denoted by the white dashed square in Fig. 1. We show the time evolution of the integrated quantities in the box in Fig. 5, such as the mass of baryonic gas, the turbulent kinetic power, and the radio power. The top row shows the projected distance between the two clusters measured from the peaks of X-ray intensity. The merger of the clump causes a rapid increase in the mass in $0.2<z<0.15$. We find that the power of solenoidal turbulence also increases around this epoch. The radio power at a fixed frequency, 140 MHz, is non-linearly dependent on the turbulent power and evolves by more than four orders of magnitude from $z=0.2$ to $z=0.1$.
@@
The radio power increases by more than two orders of magnitude in the last 1 Gyr. This rapid evolution reflects the timescale of turbulent reacceleration, which is $t_{\rm acc}\approx 500$ Myr. The 140 MHz power reaches $P_{140}\approx 4.5\times 10^{24}$ W Hz-1 at $z\approx 0.1$ when the separation of the two clusters becomes $\approx 3$ Mpc. This value is similar to the radio power of the A399–A401 radio bridge, $P_{140}\approx 1.0\times 10^{25}$ W/Hz, reported in Govoni et al. (2019).

### 08 Apr 2026 v2

# Cosmological simulation of a radio synchrotron bridge between pre-merging galaxy clusters

@@ Abstract
Methods. We perform a cosmological magneto-hydrodynamical (MHD) simulation using the Enzo code. We improve a run-time Lagrangian tracer method implemented in Enzo, and follow the trajectories of baryonic matter using $N=\mathcal{O}(10^{7})$ tracer particles. In post-processing, we conduct a parallel computation of the Fokker-Planck (FP) equation for all tracers, with cooling and reacceleration efficiencies evaluated from the local quantities recorded along each tracer trajectory.
Results. Our simulation generates a Mpc-sized radio bridge in the early stage of the cluster merger. Within a reasonable parameter range, the reacceleration model produces a broad variety of spectra. In our fiducial model, the simulated bridge matches several properties of the one found between Abell 399 and Abell 401, such as its spectral shape, intensity profile, and pixel-by-pixel correlation between radio and X-ray intensities.
@@ 1 Introduction
In the past decades, radio observations have revealed a variety of diffuse synchrotron emission in GCs, such as radio halos and cluster radio shocks (van Weeren et al. 2019). The unprecedented sensitivity of recent radio facilities, such as Low Frequency Array (LOFAR) and MeerKAT, has enabled the detection of diffuse emission on larger scales (e.g., Rajpurohit et al. 2021; Botteon et al. 2022; Cuciti et al. 2022; van Weeren et al. 2024; Rajpurohit et al. 2025). Govoni et al. (2019) reported the first discovery of a “radio bridge” at 144 MHz in A399–A401, which is diffuse emission covering the region between a pair of merging GCs.
@@
Another radio bridge has been found in the Abell 1758 system (Botteon et al. 2020), which consists of two massive galaxy clusters. However, non-detections in other similar systems indicate that such bridge-like emission is not ubiquitous in the filamentary regions between clusters (e.g., Botteon et al. 2019; Brüggen et al. 2021; Pignataro et al. 2024a). The detection of radio bridges suggests that magnetic fields and cosmic-ray electrons (CREs) can permeate into the cosmic web beyond cluster virial radii. In contrast, the non-detections indicate that specific physical conditions are required to generate a detectable radio bridge, or that the phenomenon is transient on cosmological timescales.
@@
The mechanisms responsible for magnetic field amplification and particle acceleration in the inter-cluster bridge remain uncertain. The CREs in the bridge region may be injected through the feedback processes associated with active galactic nuclei (AGNs) or star-forming galaxies (e.g., Völk et al. 1996; Berezinsky et al. 1997). However, once injected, the CREs are subjected to severe energy losses owing to synchrotron and inverse-Compton radiation, as well as Coulomb collisions. As a result, the lifetime of radio-emitting CREs ($\sim 100$ Myr) is too short for them to propagate over megaparsec scales and to produce smooth, extended bridge emission. The same problem holds for giant radio halos in GCs and is commonly referred to as the slow-diffusion problem Brunetti and Jones (2014). Reacceleration of CREs is therefore often invoked as a solution to this problem in the cluster environments (e.g., Brunetti et al. 2001; Peterson et al. 2001; Fujita et al. 2003).
@@
Numerical simulations have shown that the intra-cluster medium (ICM) and the inter-cluster bridges are permeated by shocks and turbulence driven by continuous matter accretion and episodic merger events of dark matter halos (e.g., Ryu et al. 2003; Pfrommer et al. 2006; Nelson et al. 2014; Steinwandel et al. 2024). Stochastic Fermi acceleration at these shocks or within turbulent flows may not be efficient enough to accelerate CREs directly from the thermal pool, but it can counteract radiative energy loss of the pre-existing CREs (e.g., Brunetti and Jones 2014). Secondary electrons generated through the hadronic interaction of CR protons can also contribute to diffuse synchrotron emission (e.g., Dennison 1980; Blasi and Colafrancesco 1999). However, in low-density gas regions such as cluster outskirts, the contribution from secondary electrons is expected to be smaller than in the central regions of clusters.
@@
Several previous works attempted to model energy-resolved CR distributions in GCs and LSS by combining cosmological MHD simulations with FP solvers. While earlier studies relied on MHD snapshots in post-processing (e.g., Donnert and Brunetti 2014; Pinzke et al. 2017; Winner et al. 2019; Smolinski et al. 2023; Beduzzi et al. 2024), more recent frameworks implement on-the-fly FP solvers under approximations that reduce the number of energy or momentum bins (e.g., Hopkins et al. 2022; Böss et al. 2023). We extend these approaches by incorporating the tracer-particle code Crater (Wittor et al. 2017) into the Enzo framework, and performing FP simulations in post-processing. The tracer particles generated and propagated during the runtime of Enzo better represent the baryonic flow compared to post-processing approaches such as Crater. We apply the FP solver developed in Nishiwaki et al. (2024); Nishiwaki and Asano (2025). Tests of the tracer and FP methods are presented in Appendix.
@@
Figure 1: Evolution of the projected gas density map along the Z axis in our Enzo simulation. The maps of the zoomed-in region centered on the inter-cluster filament are shown. The white square shows the (1.2 Mpc)3 box region where a radio bridge forms (Sect. 3.1).
@@ 2 Method
In Sect. 2.1, we present the set-up of the Enzo simulation of the merging clusters. During the run of the simulation, $N\approx 1.5\times 10^{7}$ tracer particles are generated to sample the baryonic fluid elements. In Sect. 2.2, we explain the methods for tracer particle generation and propagation in the Enzo simulation. The strength of solenoidal turbulence around each tracer particle is quantified using the local vorticity (Sect. 2.3). The magnetic field in our Enzo simulation may lead to an underestimation of magnetic-field amplification by the turbulent dynamo due to the limited spatial resolution as well as partially due to the use of the Dedner hyperbolic divergence-cleaning scheme to damp numerical errors in the divergence of the magnetic field. For this reason, we estimate a dynamo-amplified field in post-processing (Sect. 2.4) and adopt the larger of the simulated and modeled values. In Sect. 2.5, we introduce the FP equation adopted in this work and summarize the model parameters.
@@ 2.1 Enzo simulation
We simulate a pair of merging galaxy clusters by running the cosmological MHD code Enzo (Bryan et al. 2014). We perform a new re-simulation of the same system studied in Govoni et al. (2019) to make use of our developed tracer particle method on board Enzo. The simulation box is centered on the inter-cluster filament between merging clusters, resembling the A399–A401 system at low redshift. The root grid covers a comoving volume of $(260\penalty 10000\ \mathrm{Mpc})^{3}$, resolved with $220^{3}$ cells. Following Govoni et al. (2019), we employ the MHD solver with Dedner divergence cleaning (Dedner et al. 2002). We set a uniform comoving magnetic field with components $(B_{x},B_{y},B_{z})=(10^{-10},\,10^{-10},\,10^{-10})\penalty 10000\ \mathrm{G}$ at the initial redshift $z=30$ of the simulation. The radiative cooling of the baryonic gas is not included in this simulation.
We employ the adaptive mesh refinement (AMR) in the innermost (30 Mpc)3 region centered on the inter-cluster filament. Refined grids are introduced in the regions where the mass density of the baryonic gas or dark matter particles exceeds $0.1\rho_{\rm m}$, where $\rho_{\rm m}$ denotes the mean mass density of the Universe. The cell width is reduced by a factor of two at each refinement level. The maximum AMR level is eight, which corresponds to 4.2 comoving kpc per cell. From $z=2$, we decrease the refinement threshold to $0.01\rho_{\rm m}$ and introduce a super-Lagrangian level dependence of the refinement threshold with the exponent of 0.8-0.8. With this prescription, the refinement threshold decreases by a factor of $2^{-0.8}\simeq 0.57$ at each successive AMR level. These aggressive criteria ensures that most of the bridge volume with $\rho_{\rm gas}\gtrsim 10^{-28}$ g cm-3 is resolved with the resolution of AMR level seven (8.4 comoving kpc) in $z\lesssim 0.15$.
@@ 2.2 Lagrangian tracers
In many previous studies, the generation and propagation of Lagrangian tracer particles have been done in post-processing (e.g., Vazza et al. 2010; Wittor et al. 2017; Beduzzi et al. 2023; Smolinski et al. 2023). In this approach, particle positions are updated according to the local fluid velocity obtained by interpolating grid data from two consecutive simulation snapshots. However, because the time interval between snapshots is generally much larger than the hydrodynamical or MHD time step at high AMR levels, this method can lead to systematic over- or undershooting of tracer particle positions.
@@
We iterate over all grid cells within the $(20\penalty 10000\ \mathrm{Mpc})^{3}$ region and evaluate the baryonic mass of each cell, $M_{\rm cell}$. If $M_{\rm cell}$ lies within the range $M_{\rm trc}^{\rm min},\,M_{\rm trc}^{\rm max}]$, the cell mass is represented by a single tracer particle with $M_{\rm trc}=M_{\rm cell}$. When $M_{\rm cell}>M_{\rm trc}^{\rm max}$, which typically occurs at coarse resolution, the mass is represented by $\left\lfloor M_{\rm cell}/M_{\rm trc}^{\rm max}\right\rfloor$ tracer particles with $M_{\rm trc}=M_{\rm trc}^{\rm max}$, with one additional particle accounting for the residual mass. When $M_{\rm cell}<M_{\rm trc}^{\rm min}$, the cell mass is sampled probabilistically by generating a tracer particle with mass $M_{\rm trc}^{\rm min}$ with a probability of $M_{\rm cell}/M_{\rm trc}^{\rm min}$.
@@ 2.3 Measuring the local turbulent field
Solenoidal turbulence may play an important role in both magnetic-field amplification by the turbulent dynamo and CR reacceleration in the ICM (e.g., Cho and Vishniac 2000; Brunetti and Lazarian 2016; Vazza et al. 2018; Nishiwaki et al. 2024). Although the compressive component of turbulence may also contribute to CR acceleration (e.g., Brunetti and Lazarian 2007), we focus exclusively on the solenoidal mode in this work and defer a detailed comparison between the two modes to future studies. The local solenoidal turbulent velocity is estimated from the vorticity using the relation $\delta v(l)\approx l|\nabla\times\bm{v}|$, where $l$ denotes the spatial scale at which the vorticity is evaluated. We calculate $\nabla\times\bm{v}$ using the central-difference scheme applied to the velocity field on a stencil of $3^{3}$ cells surrounding the nearest grid cell of each tracer particle. With this approach, tracer particles probe turbulent motions on different scales, depending on the resolution of their host AMR grid. For the calculation of the dynamo-amplified magnetic field and the CR acceleration efficiency in the following sections, we normalize the turbulent velocity $\delta v$ to a reference scale of $l=L=150$ comoving kpc, assuming a Kolmogorov scaling relation $\delta v(l)\propto l^{1/3}$.
@@ 2.4 Magnetic-field amplification through turbulent dynamo
The turbulent dynamo is one of the most plausible mechanisms for the amplification of magnetic fields in galaxy clusters. Gas in the inter-cluster bridge has densities and temperatures similar to those of the cluster outskirts ($n_{e}\sim 10^{-4}$ cm-3 and $T\sim 3$ keV), so constitutes a weakly-collisional high-beta plasma. Under these conditions, microphysical instabilities can suppress the effective viscosity and allow super-Alfvénic turbulence to cascade down to small scales (e.g., Schekochihin and Cowley 2006; Brunetti and Lazarian 2007; Beresnyak and Miniati 2016; Brunetti and Jones 2014; Kunz et al. 2022). However, resolving the turbulent dynamo directly in numerical simulations is computationally very demanding. The characteristic scale of this is the so-called Alfvén scale,
@@ 2.5.3 Initial cosmic-ray distribution
Figure 5: From top to bottom, we show the time evolution of the X-ray peak distance between the clusters, baryon mass, solenoidal turbulent energy flux, and the 140 MHz radio power, and the typical equilibrium energy of CREs. The latter four quantities are calculated for the (1.2 Mpc)3 box region extracted from the inter-cluster filament (Fig. 1).
@@ 3.1 Formation of a radio bridge
To study the properties of the radio bridge, we extract a box-shaped region in the inter-cluster filament with a volume of $(1.2\penalty 10000\ {\rm Mpc})^{3}$. The region is denoted by the white dashed square in Fig. 1. We show the time evolution of the integrated quantities in the box in Fig. 5, such as the mass of baryonic gas, the turbulent kinetic power, and the radio power. The top row shows the projected distance between the two clusters measured from the peaks of X-ray intensity. The merger of the clump causes a rapid increase in the mass in $0.2<z<0.15$. We find that the power of solenoidal turbulence also increases around this epoch. The radio power at a fixed frequency, 140 MHz, is non-linearly dependent on the turbulent power and evolves by more than four orders of magnitude from $z=0.2$ to $z=0.1$.
@@
The radio power increases by more than two orders of magnitude in the last 1 Gyr. This rapid evolution reflects the timescale of turbulent reacceleration, which is $t_{\rm acc}\approx 500$ Myr. The 140 MHz power reaches $P_{140}\approx 4.5\times 10^{24}$ W Hz-1 at $z\approx 0.1$ when the separation of the two clusters becomes $\approx 3$ Mpc. This value is similar to the radio power of the A399–A401 radio bridge, $P_{140}\approx 1.0\times 10^{25}$ W Hz-1, as reported in Govoni et al. (2019).